OpenStudy (anonymous):

Simplify 3/2x + 3/x^2

OpenStudy (anonymous):



OpenStudy (anonymous):

@satellite73 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Multiply the first fraction by x/x

jimthompson5910 (jim_thompson5910):

to get \[\Large \frac{3x}{2x^2}+\frac{3}{x^2}\]

jimthompson5910 (jim_thompson5910):

now multiply the second fraction by 2/2 to get \[\Large \frac{3x}{2x^2}+\frac{6}{2x^2}\]

jimthompson5910 (jim_thompson5910):

Finally, add the fractions to get \[\Large \frac{3x+6}{2x^2}\]

OpenStudy (anonymous):

awesome thanks!!! one more?

jimthompson5910 (jim_thompson5910):

sure thing

jimthompson5910 (jim_thompson5910):

go ahead and ask

OpenStudy (anonymous):

rationalize the denominator

OpenStudy (anonymous):